OpenStudy (amistre64):

\[\int (x+1)\sqrt{2-x}\ dx\]

OpenStudy (amistre64):

im planning to distribute it thru first; or ibp it

OpenStudy (anonymous):

u=2-x and x=2-u

OpenStudy (anonymous):

\[\int\limits_{}^{}(2-u)u^{1/2}\]

OpenStudy (anonymous):

3-u instead of 2-u.

OpenStudy (anonymous):

\[\implies du = -dx\]\[-\int(3-u)\sqrt{u}\;du\]

OpenStudy (anonymous):

yes! thank you :)

OpenStudy (anonymous):

do it by parts: u = sqrt(2-x), dv = (x+1)dx; du = -1/2sqrt(2-x), v = x^2/2 + x

OpenStudy (anonymous):

eww, no need for parts here.

OpenStudy (amistre64):

wow,i had to refresh just to see all this

OpenStudy (anonymous):

you will get \[\frac{2(x-2)(x+3)\sqrt{2-x}}{5}+C\]

OpenStudy (anonymous):

Yeah don't do it by parts; do it using polpak's approach

OpenStudy (anonymous):

\[\int (u-3)\sqrt{u}\;du = \int u^\frac{3}{2}\;du - 3\int u^\frac{1}{2}du\]

OpenStudy (anonymous):

I got my stats book yesterday!

OpenStudy (anonymous):

Oh! But I have to leave now.

OpenStudy (amistre64):

\begin{array} &&&(2-x)^{1/2}\\ +&(x+1)&\frac{2}{3}(2-x)^{3/2}\\ -&1&\frac{2}{3}\frac{5}{3}(2-x)^{5/3}\\ +&0&doesnt\ matter\ here \end{array} by parts is fine, a little messy, but somethig for the teacher to wonder about lol and stats! yay, i like stats. cant do stats, but i like stats

OpenStudy (amistre64):

.... and thats spose to be 3/5 in front there

OpenStudy (amistre64):

\begin{array} &&&(2-x)^{1/2}\\ +&(x+1)&\frac{2}{3}(2-x)^{3/2}\\ -&1&\frac{2}{3}\frac{2}{5}(2-x)^{5/2}\\ +&0&doesnt\ matter\ here \end{array} \[(x+1)\frac{2}{3}(2-x)^{3/2}-\frac{2}{3}\frac{2}{5}(2-x)^{5/2}\] \[\frac{2}{3}[(x+1)(2-x)^{3/2}-\frac{2}{5}(2-x)^{5/2}]+C\] shed prolly kill me

OpenStudy (amistre64):

\[\frac{2}{3}[(x+1)(2-x)^{3/2}-\frac{2}{5}(2-x)^{5/2}]+C\] \[\frac{2(2-x)^{3/2}}{3}[(x+1)-\frac{2}{5}(2-x)]+C\] \[\frac{2(2-x)^{3/2}}{3}[\frac{5x+5-4+2x}{5}]+C\] \[\frac{2(2-x)^{3/2}}{3}[\frac{7x+1}{5}]+C\] \[\frac{2(7x+1)(2-x)^{3/2}}{15}+C\] the wolf gives me: \[\frac{-2 (2 - x)^{3/2} (3 + x)}{5}\] i wonder where i went astray

OpenStudy (amistre64):

\[\int (x+1)\sqrt{2-x}\ dx\] \[\int x(2-x)^{1/2}\ dx+\int(2-x)^{1/2}\ dx\] \[\frac{2}{3}x(2-x)^{3/2}-\frac{2}{3}\int(2-x)^{3/2}\ dx+\frac{2}{3}(2-x)^{3/2}\] \[\frac{2}{3}x(2-x)^{3/2}-\frac{2}{3}\frac{2}{5}(2-x)^{5/2}+\frac{2}{3}(2-x)^{3/2}\] but thats what i ended up with before ..... now i gotta wonder if a remember IBPs

OpenStudy (amistre64):

\[\int u\ dv=uv-\int v\ du\]

OpenStudy (amistre64):

u = x ; du = dx \[dv = (2-x)^{1/2} dx;\ v = (2/3)(2-x)^{3/2}\]

OpenStudy (anonymous):

Good luck with IBP; the integral expression will just become more complex :-(

OpenStudy (anonymous):

Don't mind me - just bookmarking this page :)

OpenStudy (amistre64):

I determined where my fauxpaux was; i forgot to pull out a -1 when i integrated my "v" choice. \begin{array}l &&\underline{\int v}\\ &\underline{du}&(2-x)^{1/2}\\ +&(x+1)&-\frac{2}{3}(2-x)^{3/2}\\ -&1&\frac{2}{3}\frac{2}{5}(2-x)^{5/2}\\ +&0&doesnt\ matter\ here \end{array} \[(x+1)-\frac{2}{3}(2-x)^{3/2}-\frac{2}{3}\frac{2}{5}(2-x)^{5/3}\] \[-\frac{2}{3}(2-x)^{3/2}\left((x+1)+\frac{2}{5}(2-x)\right)\] \[-\frac{2}{3}(2-x)^{3/2}\left(\frac{5(x+1)+2(2-x)}{5}\right)\]\[-\frac{2}{3}(2-x)^{3/2}\left(\frac{5x+5+4-2x}{5}\right)\] \[-\frac{2}{3}(2-x)^{3/2}\left(\frac{3x+9}{5}\right)\] \[-\frac{2}{\cancel{3}}(2-x)^{3/2}\frac{\cancel{3}(x+3)}{5}\] \[\frac{-2}{5}\ (2-x)^{3/2}\ (x+3)\]

OpenStudy (amistre64):

now it matches the wolfs

OpenStudy (anonymous):

that was a lot of work!

OpenStudy (anonymous):

much easier to shift a little yes?